how to make 3^x=3 into logarithmic equation. please hurry!
Answer by lwsshak3(11628)

how to make 3^x=3 into logarithmic equation. please hurry
**
3^x=3
This is the exponential form of a logarithm which states that the base(3) raised to the logarithm of the number(x)=number(3).
Logarithm form: log3(3)=x (log 3 to the base 3=x)
In this case x=1
So equation becomes log3(3)=1 (the logarithm of any base to its base is always=1)